# Capacity Bounds for Broadcast Channels with Bidirectional Conferencing Decoders

### v2

@@ IV-B Capacity Regions for Specific Channels
The above achievable rate region is derived from $\mathcal{R}_{i}^{(1)}$ given in Proposition 1 by setting $X=U=W+\bar{W}$ and $V=\varnothing$, and $\hat{Y}_{2}=Y_{2}+\hat{Z}_{2}$, where $W$ and $\bar{W}$ are two independent Gaussian variables with zero means and variances $(1-\beta)P$ and $\beta P$, respectively, and $\hat{Z}_{2}$ is a Gaussian variable (independent of all other variables) with zero mean and variance $\hat{\sigma}^{2}=1-\lambda^{2}$.
@@
Next, we present an approximate capacity result for the Gaussian BC with one-way conferencing.
@@
+$\qquad\qquad\qquad\qquad+\{C_{21}-1/2\}^{+}$ (118)
@@
Similar to Theorem 9, the above achievable rate region is derived from $\mathcal{R}_{i}^{(1)}$ given in Proposition 1 by setting $X=U=W+\bar{W}$ and $V=\varnothing$, and $\hat{Y}_{2}=Y_{2}+\hat{Z}_{2}$, where $W$ and $\bar{W}$ are two independent Gaussian variables with zero means and variances $(1-\beta)P$ and $\beta P$, respectively, and $\hat{Z}_{2}$ is a Gaussian variable (independent of all other variables) with zero mean and variance $\hat{\sigma}^{2}=1-\lambda^{2}$.
@@
The bound $\mathcal{R}_{i}^{DF-G}$ is derived from $\mathcal{R}_{i}^{(1)}$ given in Theorem 2 by setting $X=U=W+\bar{W}$ and $V=\hat{Y}_{2}=\varnothing$, where $W$ and $\bar{W}$ are two independent Gaussian variables with zero means and variances $\beta P$ and $(1-\beta)P$, respectively. Note that this inner bound is in fact derived for the channel with one-sided cooperation (i.e., it does not make use of the conferencing link $C_{21}$) using the decode-and-forward strategy alone. Nevertheless, it is a valid inner bound for the channel with bidirectional cooperation. ∎
@@
As a concluding remark for this section, the novel structure of the outer bound given in Theorem 6, in particular, the constraint (86), is crucial for deriving the exact capacity results in Theorems 7 and 8, and also the approximate capacity results in Theorems 9, 10, and 11.
@@ Appendix A Proof of Theorem 1
We only derive the constraints involving the auxiliary variable $U$, i.e., (1), (5), (6), (8), and (10). The constraints involving $V$ can be derived symmetrically. The last constraint is due to the cut-set bound.
@@
Next, we derive the constraints on the sum rate:
@@
Lastly, (11) is just the cut-set bound. By applying a standard time-sharing argument, the proof is thus complete. The cardinality bounds on $U$ and $V$ are due to Carathéodory theorem.
@@ Appendix B Proof of Proposition 1
and let $\alpha_{1}\in[0,1]$. We construct random codebooks consisting of length-$n$ codewords to transmit the common message $M_{0}$, and the private messages $M_{1}$ and $M_{2}$, which are independent and uniformly distributed over the sets $[1:2^{nR_{0}}]$, $[1:2^{nR_{1}}]$, and $[1:2^{nR_{2}}]$, respectively.
@@ B-A Encoding at the Transmitter
    If there are multiple pairs of $(b_{1},b_{2})$ that satisfy (172), let $(b_{1}^{\mathcal{T}},b_{2}^{\mathcal{T}})$ be the one with minimum $b_{1}+b_{2}$. If there is no such a pair, declare an encoding error. Based on the mutual covering lemma [29], one can easily see that for sufficiently large $n$ and small $\epsilon$, the probability of encoding error tends to zero provided that
@@ B-B Encoding and Decoding at the Receivers
    Partition the set $[1:2^{n\hat{R}_{2}}]$ into $2^{n\hat{C}_{21}}$ bins, each containing $2^{n(\hat{R}_{2}-\hat{C}_{21})}$ elements. For a given $\hat{m}_{2}\in[1:2^{n\hat{R}_{2}}]$, let $\mathfrak{C}_{21}(\hat{m}_{2})$ denote the index of the bin that $\hat{m}_{2}$ belongs to.
@@
    Given the output sequence $Y_{2}^{n}$ at the second receiver, let $\hat{m}_{2}^{\mathcal{T}}$ be the smallest integer in $[1:2^{n\hat{R}_{2}}]$ that satisfies
@@
    If there is no integer in $[1:2^{n\hat{R}_{2}}]$ that satisfies (174), then declare an error. The probability of this error tends to zero provided that
@@
    Partition the set $[1:2^{n\hat{R}_{1}}]$ into $2^{n\hat{C}_{12}^{q}}$ bins, each containing $2^{n(\hat{R}_{1}-\hat{C}_{12}^{q})}$ elements. For a given $\hat{m}_{1}\in[1:2^{n\hat{R}_{1}}]$, let $\mathfrak{C}_{12}^{q}(\hat{m}_{1})$ denote the index of the bin that $\hat{m}_{1}$ belongs to.
@@
    Given the output sequence $Y_{1}^{n}$ at the first receiver, let $\hat{m}_{1}^{\mathcal{T}}$ be the smallest integer in $[1:2^{n\hat{R}_{1}}]$ that satisfies
@@
    Note that $(m_{00}^{*},m_{11}^{*},b_{1}^{*})$ is the unique triple decoded in Step 4 according to (176)-(177). If there is no integer in $[1:2^{n\hat{R}_{1}}]$ that satisfies (182), then declare an error. The probability of this error event tends to zero provided that
@@
    Partition the set $[1:2^{nR_{00}}]$ into $2^{n\hat{C}_{12}^{d}}$ bins, each containing $2^{n(R_{00}-\hat{C}_{12}^{d})}$ elements. For a given $m_{00}\in[1:2^{nR_{00}}]$, let $\mathfrak{C}_{12}^{d}(m_{00})$ denote the index of the bin that $m_{00}$ belongs to.
@@ B-C Final Analysis
One can easily see that if $C_{21}<I(\hat{Y}_{2};Y_{2}|U,W,Y_{1})$, then the optimal $\hat{Y}_{2}=\varnothing$. Otherwise, we keep $C_{21}-I(\hat{Y}_{2};Y_{2}|U,W,Y_{1})$ as is. Then, (B-C)-(194) are equivalent to
@@
Lastly, by considering the bounds (173), (B-C)-(197), and (B-C)-(204) and by applying the Fourier–Motzkin algorithm to eliminate all the partial rates, one can derive the rate region $\mathcal{R}_{i}^{(1)}$. The proof is thus complete.
We note that the error analysis here can also be derived from the general inner bounds of [33], but the general expressions of [33] are not always easy to simplify.
@@ Appendix C Proof of Proposition 2
and let $\alpha_{2}\in[0,1]$. Similar to the proof of Proposition 1, we construct random codebooks of length-$n$ codewords to transmit the common message $M_{0}$, and the private messages $M_{1}$ and $M_{2}$, which are independent and uniformly distributed over the sets $[1:2^{nR_{0}}]$, $[1:2^{nR_{1}}]$, and $[1:2^{nR_{2}}]$, respectively.
@@
The sub-messages $M_{10}$ and $M_{20}$ are decoded at both receivers, i.e., the triple $(M_{0},M_{10},M_{20})$ is considered as a common message with rate
@@ C-A Encoding at the Transmitter
    If there are multiple pairs of $(b_{1},b_{2})$ that satisfy (213), then let $(b_{1}^{\mathcal{T}},b_{2}^{\mathcal{T}})$ be the one with minimum $b_{1}+b_{2}$. If there is no such a pair, declare encoding error. Based on the mutual covering lemma [29], one can easily see that for sufficiently large $n$ and small $\epsilon$, the probability of encoding error tends to zero provided that
@@ C-B Encoding and Decoding at the Receivers
    Partition the set $[1:2^{nR_{00}}]$ into $2^{n\hat{C}_{21}^{d}}$ bins, each containing $2^{n(R_{00}-\hat{C}_{21}^{d})}$ elements. For a given $m_{00}\in[1:2^{nR_{00}}]$, let $\mathfrak{C}_{21}^{d}(m_{00})$ denote the index of the bin that $m_{00}$ belongs to.
@@
    Partition the set $[1:2^{n\hat{R}_{2}}]$ into $2^{n\hat{C}_{21}^{q}}$ bins, each containing $2^{n(\hat{R}_{2}-\hat{C}_{21}^{q})}$ elements. For a given $\hat{m}_{2}\in[1:2^{n\hat{R}_{2}}]$, let $\mathfrak{C}_{21}^{q}(\hat{m}_{2})$ denote the index of the bin that $\hat{m}_{2}$ belongs to.
@@
    Let $\hat{m}_{2}^{\mathcal{T}}$ be the smallest integer in $[1:2^{n\hat{R}_{2}}]$ that satisfies
@@
    If there is no integer in $[1:2^{n\hat{R}_{2}}]$ that satisfies (217), then declare an error. The probability of this error tends to zero provided that
@@
    Partition the set $[1:2^{n\hat{R}_{1}}]$ into $2^{n\hat{C}_{12}}$ bins, each containing $2^{n(\hat{R}_{1}-\hat{C}_{12})}$ elements. For a given $\hat{m}_{1}\in[1:2^{n\hat{R}_{1}}]$, let $\mathfrak{C}_{12}(\hat{m}_{1})$ denote the index of the bin that $\hat{m}_{1}$ belongs to.
@@
    Given the output sequence $Y_{1}^{n}$ at the first receiver, let $\hat{m}_{1}^{\mathcal{T}}$ be the smallest integer in $[1:2^{n\hat{R}_{1}}]$ that satisfies
@@
    where $(m_{00}^{*},m_{11}^{*},b_{1}^{*})$ is the unique triple decoded in Step 6 according to (219)-(221). If there is no integer $[1:2^{n\hat{R}_{1}}]$ that satisfies (226), then declare an error. The probability of this error event tends to zero provided that
@@ C-C Final Analysis
One can easily see that if ${\alpha}_{2}C_{21}<I(\hat{Y}_{2};Y_{2}|U,W,Y_{1})$, then the optimal $\hat{Y}_{2}=\varnothing$. Thus, (232)-(235) are equivalent to
@@
Finally, by considering the bounds (237)-(238) and (243)-(244) and by applying the Fourier–Motzkin elimination algorithm, we obtain the rate region $\mathcal{R}_{i}^{(2)}$. The proof is thus complete.
@@ Appendix D Proof of Theorem 6
In other words, constraint (3) is inactive. Considering (2), (4), (7), (9), and (11) and accounting for the Markov chain $V\rightarrow X\rightarrow(Y_{1},Y_{2})$, it is easy to see that we need to maximize the following terms simultaneously:
@@
Consider first the special cases where $\lambda=\frac{b}{a}$. In this case, $Z_{1}$ and $Z_{2}$ can be thought of as being related as
@@
where $Z_{1}$, $\tilde{Z}$ are independent. Since $Y_{1}=aX+Z_{1}$, we have
@@
So, the BC is physically degraded, i.e., $X\rightarrow Y_{1}\rightarrow Y_{2}$ forms a Markov chain. The simultaneous maximization of the four mutual information terms (251)-(254) then reduces to the simultaneous maximization of $h(Y_{1}|V)$ and $h(Y_{1})$ together with minimization of $h(Y_{2}|V)$, which is achieved by jointly Gaussian $(X,V)$ as can be shown by the entropy power inequality for the degraded Gaussian broadcast channel [29].
@@
$I(X;Y_{1},Y_{2})$ $=I(X;\hat{Y},Y_{2})=I(X;\hat{Y}).$ (266)
Now, utilizing (265)-(266) for bounding the maximum values of (251)-(254), we see that after expressing the mutual information in terms of differential entropy, the optimization is effectively to maximize $h(\hat{Y}|V)$, $h(Y_{1}|V)$ and $h(\hat{Y})$, and simultaneously to minimize $h(Y_{2}|V)$.
@@
Considering (272), (278), and (284) together with the fact that a Gaussian $X\sim\mathcal{N}(0,P)$ maximizes $h(\hat{Y})$ allows us to conclude that a jointly Gaussian $(X,V)$ simultaneously maximizes the mutual information terms in (251)-(254) and (265)-(266).

### v4

@@ IV-B Capacity Regions for Specific Channels
The above achievable rate region is derived from $\mathcal{R}_{i}^{(1)}$ given in Proposition 1 by setting $X=U=W+\bar{W}$ and $V=\hat{Y}_{1}=\varnothing$, and $\hat{Y}_{2}=Y_{2}+\hat{Z}_{2}$, where $W$ and $\bar{W}$ are two independent Gaussian variables with zero means and variances $(1-\beta)P$ and $\beta P$, respectively, and $\hat{Z}_{2}$ is a Gaussian variable (independent of all other variables) with zero mean and variance $\hat{\sigma}^{2}=1-\lambda^{2}$.
@@
Next, we present an approximate capacity result for the Gaussian BC with one-sided conferencing link.
@@
$\qquad\qquad\qquad\qquad+\{C_{21}-1/2\}^{+}$ (118)
@@
Similar to Theorem 9, the above achievable rate region is derived from $\mathcal{R}_{i}^{(1)}$ given in Proposition 1 by setting $X=U=W+\bar{W}$ and $V=\hat{Y}_{1}=\varnothing$, and $\hat{Y}_{2}=Y_{2}+\hat{Z}_{2}$, where $W$ and $\bar{W}$ are two independent Gaussian variables with zero means and variances $(1-\beta)P$ and $\beta P$, respectively, and $\hat{Z}_{2}$ is a Gaussian variable (independent of all other variables) with zero mean and variance $\hat{\sigma}^{2}=1-\lambda^{2}$.
@@
The bound $\mathcal{R}_{i}^{DF-G}$ is derived from $\mathcal{R}_{i}^{(1)}$ given in Theorem 2 by setting $X=U=W+\bar{W}$ and $V=\hat{Y}_{1}=\hat{Y}_{2}=\varnothing$, where $W$ and $\bar{W}$ are two independent Gaussian variables with zero means and variances $\beta P$ and $(1-\beta)P$, respectively. Note that this inner bound is in fact derived for the channel with one-sided cooperation (i.e., it does not make use of the conferencing link $C_{21}$) using the decode-and-forward strategy alone. Nevertheless, it is a valid inner bound for the channel with bidirectional cooperation. ∎
@@
As a concluding remark for this section, the novel structure of the outer bound given in Theorem 6, in particular, the constraint (86), is crucial for deriving the exact capacity results in Theorems 7 and 8, and also the approximate capacity results in Theorems 9, 10, and $1$.
@@ Appendix A Proof of Theorem 1
We only derive the constraints that include the auxiliary variable $U$, i.e., (1), (5), (6), (8), and (10). The constraints that include $V$ can be derived symmetrically. The last constraint (11) is due to the cut-set bound.
@@
Next, we derive the constraints on the sum rate, starting from Fano’s inequality:
@@
By applying a standard time-sharing argument, the proof is thus complete. The cardinality bounds on $U$ and $V$ are due to Carathéodory theorem [29].
@@ Appendix B Proof of Proposition 1
and let $\alpha_{1}\in[0,1]$. We construct random codebooks consisting of length-$n$ codewords to transmit the common message $M_{0}$, and the private messages $M_{1}$ and $M_{2}$, which are independent and uniformly distributed over the sets $1:2^{nR_{0}}]$, $1:2^{nR_{1}}]$, and $1:2^{nR_{2}}]$, respectively.
@@ B-A Encoding at the Transmitter
    If there are multiple pairs of $(b_{1},b_{2})$ that satisfy (172), let $(b_{1}^{\mathcal{T}},b_{2}^{\mathcal{T}})$ be the one with minimum $b_{1}$. If multiple such pairs have the same minimum $b_{1}$, choose the one with the minimum $b_{2}$. If no pair $(b_{1},b_{2})$ satisfies (172), declare an encoding error. Based on the mutual covering lemma [29], one can easily see that for sufficiently large $n$ and small $\epsilon$, the probability of encoding error tends to zero provided that
@@ B-B Encoding and Decoding at the Receivers
    Partition the set $1:2^{n\hat{R}_{2}}]$ into $2^{n\hat{C}_{21}}$ bins, each containing $2^{n(\hat{R}_{2}-\hat{C}_{21})}$ elements. For a given $\hat{m}_{2}\in[1:2^{n\hat{R}_{2}}]$, let $\mathfrak{C}_{21}(\hat{m}_{2})$ denote the index of the bin that $\hat{m}_{2}$ belongs to.
@@
    Given the output sequence $Y_{2}^{n}$ at the second receiver, let $\hat{m}_{2}^{\mathcal{T}}$ be the smallest integer in $1:2^{n\hat{R}_{2}}]$ that satisfies
@@
    If there is no integer in $1:2^{n\hat{R}_{2}}]$ that satisfies (174), then declare an error. The probability of this error tends to zero provided that
@@
    Partition the set $1:2^{n\hat{R}_{1}}]$ into $2^{n\hat{C}_{12}^{q}}$ bins, each containing $2^{n(\hat{R}_{1}-\hat{C}_{12}^{q})}$ elements. For a given $\hat{m}_{1}\in[1:2^{n\hat{R}_{1}}]$, let $\mathfrak{C}_{12}^{q}(\hat{m}_{1})$ denote the index of the bin that $\hat{m}_{1}$ belongs to.
@@
    Given the output sequence $Y_{1}^{n}$ at the first receiver, let $\hat{m}_{1}^{\mathcal{T}}$ be the smallest integer in $1:2^{n\hat{R}_{1}}]$ that satisfies
@@
    Note that $(m_{00}^{*},m_{11}^{*},b_{1}^{*})$ is the unique triple decoded in Step 4 according to (176)-(177). If there is no integer in $1:2^{n\hat{R}_{1}}]$ that satisfies (182), then declare an error. The probability of this error event tends to zero provided that
@@
    Partition the set $1:2^{nR_{00}}]$ into $2^{n\hat{C}_{12}^{d}}$ bins, each containing $2^{n(R_{00}-\hat{C}_{12}^{d})}$ elements. For a given $m_{00}\in[1:2^{nR_{00}}]$, let $\mathfrak{C}_{12}^{d}(m_{00})$ denote the index of the bin that $m_{00}$ belongs to.
@@ B-C Final Analysis
One can easily see that if $C_{21}<I(\hat{Y}_{2};Y_{2}|U,W,Y_{1})$, then the optimal $\hat{Y}_{2}=\varnothing$. Otherwise, we keep $C_{21}-I(\hat{Y}_{2};Y_{2}|U,W,Y_{1})$ as is.
Then, (B-C)-(194) are equivalent to
@@
Lastly, considering the bounds in (173), (B-C)-(197), and (B-C)-(204) together, by applying the Fourier–Motzkin algorithm to eliminate all the partial rates, one can derive the rate region $\mathcal{R}_{i}^{(1)}$. The proof is thus complete.
We note that the error probability analysis here can also be derived from the general approach of [34], but the general expressions in [34] are not always easy to simplify.
@@ Appendix C Proof of Proposition 2
and let $\alpha_{2}\in[0,1]$. Similar to the proof of Proposition 1, we construct random codebooks of length-$n$ codewords to transmit the common message $M_{0}$, and the private messages $M_{1}$ and $M_{2}$, which are independent and uniformly distributed over the sets $1:2^{nR_{0}}]$, $1:2^{nR_{1}}]$, and $1:2^{nR_{2}}]$, respectively.
@@ C-A Encoding at the Transmitter
    If there are multiple pairs of $(b_{1},b_{2})$ that satisfy (213), let $(b_{1}^{\mathcal{T}},b_{2}^{\mathcal{T}})$ be the one with minimum $b_{1}$. If multiple such pairs have the same minimum $b_{1}$, choose the one with the minimum $b_{2}$. If no pair $(b_{1},b_{2})$ satisfies (213), declare an encoding error. Based on the mutual covering lemma [29], one can easily see that for sufficiently large $n$ and small $\epsilon$, the probability of encoding error tends to zero provided that
@@ C-B Encoding and Decoding at the Receivers
    Partition the set $1:2^{nR_{00}}]$ into $2^{n\hat{C}_{21}^{d}}$ bins, each containing $2^{n(R_{00}-\hat{C}_{21}^{d})}$ elements. For a given $m_{00}\in[1:2^{nR_{00}}]$, let $\mathfrak{C}_{21}^{d}(m_{00})$ denote the index of the bin that $m_{00}$ belongs to.
@@
    Partition the set $1:2^{n\hat{R}_{2}}]$ into $2^{n\hat{C}_{21}^{q}}$ bins, each containing $2^{n(\hat{R}_{2}-\hat{C}_{21}^{q})}$ elements. For a given $\hat{m}_{2}\in[1:2^{n\hat{R}_{2}}]$, let $\mathfrak{C}_{21}^{q}(\hat{m}_{2})$ denote the index of the bin that $\hat{m}_{2}$ belongs to.
@@
    Let $\hat{m}_{2}^{\mathcal{T}}$ be the smallest integer in $1:2^{n\hat{R}_{2}}]$ that satisfies
@@
    If there is no integer in $1:2^{n\hat{R}_{2}}]$ that satisfies (217), then declare an error. The probability of this error tends to zero provided that
@@
    Partition the set $1:2^{n\hat{R}_{1}}]$ into $2^{n\hat{C}_{12}}$ bins, each containing $2^{n(\hat{R}_{1}-\hat{C}_{12})}$ elements. For a given $\hat{m}_{1}\in[1:2^{n\hat{R}_{1}}]$, let $\mathfrak{C}_{12}(\hat{m}_{1})$ denote the index of the bin that $\hat{m}_{1}$ belongs to.
@@
    Given the output sequence $Y_{1}^{n}$ at the first receiver, let $\hat{m}_{1}^{\mathcal{T}}$ be the smallest integer in $1:2^{n\hat{R}_{1}}]$ that satisfies
@@
    where $(m_{00}^{*},m_{11}^{*},b_{1}^{*})$ is the unique triple decoded in Step 6 according to (219)-(221). If there is no integer $1:2^{n\hat{R}_{1}}]$ that satisfies (226), then declare an error. The probability of this error event tends to zero provided that
@@ C-C Final Analysis
One can easily see that if ${\alpha}_{2}C_{21}<I(\hat{Y}_{2};Y_{2}|U,W,Y_{1})$, then the optimal $\hat{Y}_{2}=\varnothing$.
Thus, (232)-(235) are equivalent to
@@
Finally, considering the bounds (237)-(238) and (243)-(244) together, by applying the Fourier–Motzkin algorithm to eliminate all the partial rates, we obtain the rate region $\mathcal{R}_{i}^{(2)}$. The proof is thus complete.
@@ Appendix D Proof of Theorem 6
In other words, constraint (3) is inactive.
Considering (2), (4), (7), (9), and (11) and accounting for the Markov chain $V\rightarrow X\rightarrow(Y_{1},Y_{2})$, it is easy to see that we need to maximize the following terms simultaneously:
@@
Consider first the special case where $\lambda=\frac{b}{a}$. In this case, $Z_{1}$ and $Z_{2}$ can be thought of as being related as
@@
where $_{1}, Z~\tilde{Z}$ are independent. Since $Y_{1}=aX+Z_{1}$, we have
@@
So, the BC is physically degraded (because $|a|\geq|b|$), i.e., $X\rightarrow Y_{1}\rightarrow Y_{2}$ forms a Markov chain.
In this special case, the simultaneous maximization of the four mutual information terms (251)-(254) then reduces to the simultaneous maximization of $h(Y_{1}|V)$ and $h(Y_{1})$ together with minimization of $h(Y_{2}|V)$, which is achieved by jointly Gaussian $(X,V)$ as can be shown by the entropy power inequality in the proof of the capacity region of the degraded Gaussian broadcast channel [29].
@@
[…]
Now, utilizing (265)-(266) for bounding the maximum values of (251)-(254), we see that after expressing the mutual information in terms of differential entropy, the optimization is effectively to maximize $h(\hat{Y}|V)$, $h(Y_{1}|V)$ and h(Y^)h($\hat${Y}), and simultaneously to minimize $h(Y_{2}|V)$.
@@
Considering (272), (278), and (284) together with the fact that a Gaussian $X\sim\mathcal{N}(0,P)$ maximizes h(Y^)h($\hat${Y}) allows us to conclude that a jointly Gaussian $(X,V)$ simultaneously maximizes the mutual information terms in (251)-(254) and (265)-(266).